can you please explain how you get the answer step by step
Mr. Wise bought $1950 woth of stock, some at $3 per share and some at $4.50 per share. If he bought a total of 450 shares of stock, how many of each did he buy?

thank you
april

Answer by nerdybill(7384)

Mr. Wise bought $1950 woth of stock, some at $3 per share and some at $4.50 per share. If he bought a total of 450 shares of stock, how many of each did he buy?
.
Let x = number of $3.50 shares purchased
and y = number of $4.50 shares purchased
.
From: "Mr. Wise bought $1950 woth of stock" we get
3.50x + 4.50y = 1950 (equation 1)
.
From: "If he bought a total of 450 shares of stock" we get
x + y = 450 (equation 2)
.
Apply the "addition" method:
3.50x + 4.50y = 1950
x + y = 450
.
Multiply the bottom equation by -3.50:
3.50x + 4.50y = 1950
-3.50x - 3.50y = -1575
.
Add the two equations together:
3.50x + 4.50y = 1950
-3.50x - 3.50y = -1575
--------------------------
1.50y = 375
y = 250 shares at $4.50
.
Shares at $3.50, substitute above into equation 2:
x + y = 450
x + 250 = 450
x = 200 shares at $3.50